# Robust Deep Reinforcement Learning against Adversarial Behavior Manipulation

### v2

@@ F.2 Experiments on MiniGrid environments
Results in coordinate states
@@
Results in vision-based states
@@ H.1 Full results in defense performance evaluation
In our evaluation of defense methods in Section 7, we only provide the results of the best attack. In this section, we present the results of all attacks in 8. Targeted PGD is ineffective against robustly trained victims. We observed a trend where the Reward Maximization Attack tended to be slightly more effective than BIA when attacking highly robust victims. When the victim’s robustness is high, it becomes difficult to make the victim behave like the target policy, which may cause the learning process in BIA to fail or collapse. On the other hand, in the Reward Maximization Attack, the reward serves as a good guidepost, allowing learning to proceed even when the victim’s robustness is high.
@@ H.2 Impact of Regularization Coefficient
We analyze how policy smoothing affects both robustness and original performance. The strength of the regularization is determined by the coefficient $\gamma$. Accordingly, we evaluate the clean rewards and best attack rewards for SA-PPO and TDRT-PPO with $\gamma\in\{0.03,0.1,0.2,0.3,0.5\}$. The results are shown in Table 9, and all other settings follow those in Section 7 (where $\gamma$=0.3).
Robustness
$\gamma$=0.5, the best attack rewards remain unchanged compared to $\gamma$=0.3 for both TDRT-PPO and SA-PPO, indicating that $\gamma$=0.3 provides sufficient regularization, and that further increasing it does not enhance robustness. Conversely, as $\gamma$ decreases, the best attack rewards decline. Notably, TDRT-PPO fails to maintain robustness when $\gamma$=0.03, highlighting the need for a certain level of regularization to achieve robust performance.
Original Performance
Both TDRT-PPO and SA-PPO experience a slight drop in performance as $\gamma$ increases, likely because stronger regularization reduces the expressiveness of the policy. For SA-PPO, lowering $\gamma$ improves performance, suggesting that $\gamma$=0.3 may be overly stringent and that reducing it helps recover the policy’s representational capacity.
@@
Table 9: Comparison between TDRT-PPO and SA-PPO with different $\gamma$ values. Each value is the average episode rewards $\pm$ standard deviation over 50 episodes. Clean Rewards are the rewards for the victim’s tasks (no attack). Best attack rewards represent the highest attack reward among the five types of adversarial attacks. The attack budget is set to $\epsilon=0.3$. In SA-PPO, which does not apply time discounting, ensuring sufficient robustness leads to a significant drop in performance on the original task. In contrast, TDRT-PPO, which applies time discounting, achieves high robustness while preserving the original-task performance.
Adv Task $\gamma$ TDRT-PPO (ours) SA-PPO
@@
H.3 Traing time efficiency
@@ Appendix I Implementation Details
In this section, we provide a detailed explanation of the implementation. Our code is available at https://anonymous.4open.science/r/Behavior-Targeted-Attack-6643
@@ I.1 Environment Details
We conducted the experiments described in Section 7 using Meta-World[34], a benchmark that simulates robotic arm manipulation. All tasks in Meta-World share a common 39-dimensional continuous state space and a 4-dimensional continuous action space. In our experiments, we use four tasks: window-close, window-open, drawer-close, drawer-open, faucet-close, faucet-open, handle-press-sid, handle-pull-side, door-lock, and door-unlock. The objective of each task is to move a specific object to a designated position.
@@
The reward design is specific to each task in MetaWorld. For example, the reward functions for window-close and window-open tasks are designed independently. Therefore, when attacking a victim trained on the window-close task to perform a window-open task, this differs from an untargeted attack since it does not simply minimize the victim’s reward.
@@ I.2 Defense baseline details
[7]: This method aims to increase the smoothness of the policy’s action outputs by a regularizer. The difference between TDRT-PPO and SA-PPO is that SA-PPO does not apply time discounting in the regularization.
@@
[11]: ATLA-PPO is an adversarial training method that alternates between training a victim and an adversary. The adversary is trained to generate perturbations that produce the worst-case cumulative reward for the victim by leveraging the SA-MDP framework. The agent is then trained to optimize its policy against this strong adversary, resulting in improved robustness to adversarial attacks.
@@
[8]: PA-ATLA-PPO extends ATLA-PPO by integrating the Policy Adversarial Actor Director (PA-AD) framework, which improves adversarial attack generation. PA-AD utilizes a "director" to determine optimal policy perturbation directions and an "actor" to generate corresponding state perturbations, ensuring efficient and theoretically optimal adversarial attacks. Unlike ATLA-PPO, which trains an adversary directly using reinforcement learning, PA-ATLA-PPO’s decoupled approach enhances efficiency and scalability in environments with large state spaces.
(iii) WocaR-PPO
[24]: WocaR-PPO trains the policy to maximize the worst-case cumulative reward under adversarial attacks. Unlike adversarial training methods such as ATLA-PPO, which involve learning alongside an adversary, WocaR-PPO uses a computationally feasible approach to estimate the worst-case cumulative reward without requiring additional interaction with the environment. Additionally, regularization is applied to improve the smoothness of the policy, focusing specifically on critical states where significant reward drops are likely to occur based on state importance weights.
(iv) RAD-PPO
[32]: RAD-PPO is a regret-based defense approach. Regret represents the difference between the value without an attack and the value under an attack. RAD-PPO aims to achieve robustness against adversarial attacks by learning a policy that minimizes regret at each step. Since regret is defined based on the rewards obtained by the victim, regret-based defense methods primarily assume untargeted adversaries. Therefore, the robustness of this approach cannot be fully guaranteed against the behavior-targeted attack. It is worth noting that the implementation code for RAD-PPO is not publicly available, so we implemented it ourselves based on the details provided in the paper.
@@ I.3 Hyperparameter details
In this section, we discuss the hyperparameters used in our experiments. In general, our hyperparameter settings follow [11].
@@ I.4.1 Implementation details of the targeted PGD attack
3:  $\epsilon_{\text{step}}\leftarrow\epsilon/T$
4:  $s_{\text{min}}\leftarrow s-\epsilon$, $s_{\text{max}}\leftarrow s+\epsilon$
5:  $\hat{s}\leftarrow s+\text{Uniform}(-\epsilon_{\text{step}},\epsilon_{\text{step}})$
6:  $a_{\text{tgt}}\sim\pi_{\text{tgt}}(\hat{s}|\cdot)$

### v3

@@ F.2 Experiments on MiniGrid environments
Results in coordinate states.
@@
Results in vision-based states.
@@
Defense Performance Comparison.
Table 8: Comparison of defense performances under the best attack in MinGrid environments. Each value represents the average episode reward $\pm$ standard deviation over 50 episodes. Higher clean rewards and lower best attack rewards indicate better defense performance.
Task Method Clean Reward ($\uparrow$) Best Attack Reward ($\uparrow$)
$8\times 8$ PPO 0.93 $\pm$ 0.04 0.12 $\pm$ 0.32
SA-PPO (smoothing w/o time-discounting) 0.93 $\pm$ 0.05 0.78 $\pm$ 0.11
TDRT-PPO (ours, smoothing w/ time-discounting) 0.92 $\pm$ 0.08 0.74 $\pm$ 0.15
$16\times 16$ PPO 0.92 $\pm$ 0.03 0.24 $\pm$ 0.22
SA-PPO (smoothing w/o time-discounting) 0.92 $\pm$ 0.01 0.71 $\pm$ 0.12
TDRT-PPO (ours, smoothing w/ time-discounting) 0.93 $\pm$ 0.04 0.74 $\pm$ 0.04
Analogously to the MuJoCo experiments, we compare TDRT with the baseline method in the MiniGrid environment. Because attack performance is insufficient when using vision-based observations, we restrict our evaluation to coordinate-state settings, while keeping all other experimental configurations identical to the MuJoCo setup. The results are summarized in Table 8. TDRT-PPO and SA-PPO exhibit similar robustness and clean performance, which we conjecture is because MiniGrid is a relatively easy task in which both methods nearly saturate the achievable performance, leaving little room for noticeable differences in clean performance.
@@ G.2 Attack budget efficiency
G.4 Evaluating Attack Performance with ASR
Table 10: Comparison of attack performances. Each value represents the average attack success rate (ASR) over 50 episodes. Each parenthesis indicates (Access model, Adversary’s knowledge). We set the attack budget $\epsilon=0.3$.
Task
Targeted PGD
(white-box)
Rew Max (PA-AD)
(black-box)
Rew Max (SA-RL)
(black-box)
BIA-ILfD (ours)
(black-box)
BIA-ILfO (ours)
(no-box)
Random
(no-box)
window-close 0.08 0.82 1.00 0.72 0.74 0.00
window-open 0.08 0.02 0.06 0.12 0.06 0.00
drawer-close 0.68 0.68 1.00 1.00 1.00 0.22
drawer-open 0.00 0.00 0.00 0.00 0.00 0.00
faucet-close 0.00 0.04 0.68 0.68 0.60 0.00
faucet-open 0.00 0.00 0.00 0.12 0.08 0.00
handle-press-side 0.16 1.00 0.98 1.00 1.00 0.02
handle-pull-side 0.06 0.70 0.92 1.00 1.00 0.00
door-lock 0.04 0.04 0.46 0.50 0.42 0.00
door-unlock 0.00 0.62 0.68 0.64 0.62 0.00
In this section, we evaluate the performance of the attack methods using the attack success rate (ASR). Success is determined by the task success flag provided by the Meta-World environment. For each setting, we run 50 episodes and report ASR as the fraction of episodes marked as successful. All other experimental settings follow the main text.
The results are shown in Table 10. Consistent with the evaluation based on episode reward, the results indicate that BIA is effective, supporting that the attack performance of our proposed method generalizes across evaluation metrics.
G.5 Computational Cost of Training Adversarial Policies
In this section, we compare the computational cost of different adversarial attack methods. We divide computational cost into two components: training-time cost and test-time cost. This distinction is important because different attack families place the computational burden at different phases. For example, BIA requires training an adversarial policy, but its test-time execution is simply a single forward pass. In contrast, Targeted PGD requires no training, but incurs optimization overhead at every timestep during test-time rollout. Table 11 summarizes the computational cost measured on a single NVIDIA H100 GPU.
Table 11: Comparison of computational cost.
Targeted PGD Rew Max (PA-AD) Rew Max (SA-RL) BIA-ILfD (ours) BIA-ILfO (ours)
Training time 3.6 h 4.3 h 6.2 h 6.1 h
Test time 37 sec / episode
BIA vs. reward-based white-box attacks (Rew Max).
BIA has a higher training cost than Rew Max (PA-AD). This is mainly because imitation learning–based BIA must train both a policy and a discriminator, while Rew Max only optimizes the adversarial policy. On the other hand, BIA only requires demonstrations of the target behavior, whereas Rew Max assumes access to the target reward function, which is often harder to obtain in practice.
BIA vs. Targeted PGD.
As described above, Targeted PGD has no training-time cost, but incurs a test-time optimization cost. When many episodes need to be attacked (e.g., long deployments or large-scale evaluation), BIA’s training cost is spread out over many episodes, and BIA becomes more computationally efficient overall than repeatedly running Targeted PGD.
@@ H.1 Full results in defense performance evaluation
In our evaluation of defense methods in Section 7, we only provide the results of the best attack. In this section, we present the results of all attacks in Table 12. Targeted PGD is ineffective against robustly trained victims. We observed a trend where the Reward Maximization Attack tended to be slightly more effective than BIA when attacking highly robust victims. When the victim’s robustness is high, it becomes difficult to make the victim behave like the target policy, which may cause the learning process in BIA to fail or collapse. On the other hand, in the Reward Maximization Attack, the reward serves as a good guidepost, allowing learning to proceed even when the victim’s robustness is high.
@@ H.2 Impact of Regularization Coefficient
We analyze how policy smoothing affects both robustness and original performance. The strength of the regularization is determined by the coefficient $\lambda$. Accordingly, we evaluate the clean rewards and best attack rewards for SA-PPO and TDRT-PPO with $\lambda\in\{0.03,0.1,0.2,0.3,0.5\}$. The results are shown in Table 13, and all other settings follow those in Section 7 (where $\lambda$=0.3).
Robustness.
For $\lambda$=0.5, the best attack rewards remain unchanged compared to $\gamma$=0.3 for both TDRT-PPO and SA-PPO, indicating that $\lambda$=0.3 provides sufficient regularization, and that further increasing it does not enhance robustness. Conversely, as $\lambda$ decreases, the best attack rewards decline. Notably, TDRT-PPO fails to maintain robustness when $\lambda$=0.03, highlighting the need for a certain level of regularization to achieve robust performance.
Original Performance.
Both TDRT-PPO and SA-PPO experience a slight drop in performance as $\lambda$ increases, likely because stronger regularization reduces the expressiveness of the policy. For SA-PPO, lowering $\lambda$ improves performance, suggesting that $\lambda$=0.3 may be overly stringent and that reducing it helps recover the policy’s representational capacity.
@@
Table 13: Comparison between TDRT-PPO and SA-PPO with different $\lambda$ values. Each value is the average episode rewards $\pm$ standard deviation over 50 episodes. Clean Rewards are the rewards for the victim’s tasks (no attack). Best attack rewards represent the highest attack reward among the five types of adversarial attacks. The attack budget is set to $\epsilon=0.3$. In SA-PPO, which does not apply time discounting, ensuring sufficient robustness leads to a significant drop in performance on the original task. In contrast, TDRT-PPO, which applies time discounting, achieves high robustness while preserving the original-task performance.
Adv Task $\lambda$ TDRT-PPO (ours) SA-PPO
@@
H.3 Training time efficiency
@@ Appendix I Implementation Details
In this section, we provide a detailed explanation of the implementation.
@@ I.1 Environment Details
We conducted the experiments described in Section 7 using Meta-World(Yu et al., 2020), a benchmark that simulates robotic arm manipulation. All tasks in Meta-World share a common 39-dimensional continuous state space and a 4-dimensional continuous action space. In our experiments, we use four tasks: window-close, window-open, drawer-close, drawer-open, faucet-close, faucet-open, handle-press-side, handle-pull-side, door-lock, and door-unlock. The objective of each task is to move a specific object to a designated position.
@@
The reward design is specific to each task in Meta-World. For example, the reward functions for window-close and window-open tasks are designed independently. Therefore, when attacking a victim trained on the window-close task to perform a window-open task, this differs from an untargeted attack since it does not simply minimize the victim’s reward.
@@ I.2 Defense baseline details
(Zhang et al., 2020b): This method aims to increase the smoothness of the policy’s action outputs by a regularizer. The difference between TDRT-PPO and SA-PPO is that SA-PPO does not apply time discounting in the regularization.
@@
(Zhang et al., 2021): ATLA-PPO is an adversarial training method that alternates between training a victim and an adversary. The adversary is trained to generate perturbations that produce the worst-case cumulative reward for the victim by leveraging the SA-MDP framework. The agent is then trained to optimize its policy against this strong adversary, resulting in improved robustness to adversarial attacks.
@@
(Sun et al., 2022): PA-ATLA-PPO extends ATLA-PPO by integrating the Policy Adversarial Actor Director (PA-AD) framework, which improves adversarial attack generation. PA-AD utilizes a "director" to determine optimal policy perturbation directions and an "actor" to generate corresponding state perturbations, ensuring efficient and theoretically optimal adversarial attacks. Unlike ATLA-PPO, which trains an adversary directly using reinforcement learning, PA-ATLA-PPO’s decoupled approach enhances efficiency and scalability in environments with large state spaces.
(iv) WocaR-PPO
(Liang et al., 2022): WocaR-PPO trains the policy to maximize the worst-case cumulative reward under adversarial attacks. Unlike adversarial training methods such as ATLA-PPO, which involve learning alongside an adversary, WocaR-PPO uses a computationally feasible approach to estimate the worst-case cumulative reward without requiring additional interaction with the environment. Additionally, regularization is applied to improve the smoothness of the policy, focusing specifically on critical states where significant reward drops are likely to occur based on state importance weights.
(v) RAD-PPO
(Belaire et al., 2024): RAD-PPO is a regret-based defense approach. Regret represents the difference between the value without an attack and the value under an attack. RAD-PPO aims to achieve robustness against adversarial attacks by learning a policy that minimizes regret at each step. Since regret is defined based on the rewards obtained by the victim, regret-based defense methods primarily assume untargeted adversaries. Therefore, the robustness of this approach cannot be fully guaranteed against the behavior-targeted attack. It is worth noting that the implementation code for RAD-PPO is not publicly available, so we implemented it ourselves based on the details provided in the paper.
@@ I.3 Hyperparameter details
In this section, we discuss the hyperparameters used in our experiments. In general, our hyperparameter settings follow (Zhang et al., 2021).
@@ I.4.1 Implementation details of the targeted PGD attack
[…]